# Revisiting Graph Neural Networks on Graph-level Tasks: Comprehensive Experiments, Analysis, and Improvements

### v1

@@ 4. Experiments
We conduct experiments to analyze baselines and our model on both graph classification and regression tasks. Specifically, we introduce datasets and baselines in Section 4.1 and Section 4.2. We present the hyperparameter setting and the evaluation metric in Section 4.3 and Section 4.4, respectively. Effectiveness and efficiency are discussed in Section 4.5 and Section 4.6. Additionally, we evaluate baselines on more scenarios in Section 4.7.
@@ 4.1. Datasets
We use graphs from four domains, including social networks (SN), biology (BIO), chemistry (CHE), and geometric (GE) graphs. In particular, for dataset splitting, if a standard split is publicly available and commonly used in existing works, we adopt the standard split. If there is no standard split, we use a 10-fold cross-validation approach. The detailed data statistics and split ratio are in Table 2.
@@ 4.1.1. Social Networks
  • IMDB-B (Morris et al., 2020) and IMDB-M (Morris et al., 2020). Two movie datasets from the IMDB (imd, 2024), where nodes are actors, edges denote co-appearances, and labels indicate movie genres. IMDB-B and IMDB-M are binary and multiple-class classification datasets, respectively.
  • REDDIT-B (Morris et al., 2020). A binary social network dataset from Reddit (red, 2024), where nodes represent users, edges indicate interactions, and graphs are labeled as QA-based or discussion-based.
  • COLLAB-M (Morris et al., 2020). A multi-class scientific collaboration network, where nodes are researchers and edges indicate co-authorship. Each graph is labeled with different research fields.
@@ 4.1.2. Biology
  • PROTEIN-B (Morris et al., 2020). A binary protein dataset where nodes represent secondary structure elements, edges denote proximity, and graphs indicate if the molecule is an enzyme.
  • DD-B (Morris et al., 2020). Derived from a molecular protein dataset, where nodes represent amino acids, and edges indicate adjacency. Each graph is labeled to classify different types of proteins.
  • ENZYMES (Morris et al., 2020). Derived from the BRENDA enzyme database, it represents nodes as secondary structural elements, edges as adjacency, and labels each graph with one of six EC classes.
  • MolHIV (Hu et al., 2020b) and MolTox21 (Hu et al., 2020b). They are molecular datasets, where nodes represent atoms, and edges are chemical bonds. Labels in MolHIV indicate whether the molecule has HIV inhibition, and labels in MolTox21 indicate the toxicity type.
@@ 4.1.3. Chemistry
MUTAG, NCI1, MolBACE, and MolPCBA are molecular graph datasets, with nodes represent atoms, node labels indicating the atomic species (e.g., N, C, O), and edges represent chemical bonds between atoms.
  • MUTAG (Morris et al., 2020). The labels are whether a molecule has mutagenic activity, referring to its potential to cause genetic mutations.
  • NCI1 (Morris et al., 2020). The labels specify whether a molecule shows a positive response in assays for specific anticancer activity.
  • MolBACE (Hu et al., 2020b). The labels indicate whether a molecule inhibits the BACE enzyme (Beta-secretase 1).
  • MolPCBA (Hu et al., 2020b). The labels represent the activity of a molecule across 128 different bioassays, indicating whether the compound is active in a given experimental condition.
Table 2. Data statistics. CV denotes Cross-Validation.
Type Data Graphs $\overline{\mbox{Nodes}}$ $\overline{\mbox{Edges}}$ Feat Classes Split Ratio
SN IMDB-B 1,000 19.77 96.53 N/A 2 10-fold CV
IMDB-M 1,500 13.00 65.94 N/A 3 10-fold CV
REDDIT-B 2,000 429.63 497.75 N/A 2 10-fold CV
COLLAB 5,000 74.49 2,457.78 N/A 3 10-fold CV
BIO PROTEINS 1,113 39.06 72.82 3 2 10-fold CV
DD 1,178 284.32 715.66 89 2 10-fold CV
ENZYMES 600 32.63 62.14 3 6 10-fold CV
MolHIV 41,127 25.50 54.93 9 2 8/1/1
MolTox21 7,831 18.90 39.22 9 12 8/1/1
CHE MUTAG 188 17.93 19.79 7 2 10-fold CV
NCI1 4,110 29.87 32.30 37 2 10-fold CV
MolBACE 1,513 34.13 36.91 204 2 8/1/1
MolPCBA 437,929 26.00 28.10 128 128 8/1/1
GE $\{\mathbf{3,4,5}\}$-Cycle 5,000 18.80 31.34 1 1 3/2/5
$\mathbf{\{6,7,8\}}$-Cycle 5,000 18.80 31.34 1 1 3/2/5
$\mathbf{\{4,5,6\}}$-Path 5,000 18.80 31.34 1 1 3/2/5
4-Clique 5,000 18.80 31.34 1 1 3/2/5
Tailed Tri. 5,000 18.80 31.34 1 1 3/2/5
Chordal Cyc. 5,000 18.80 31.34 1 1 3/2/5
Triangle Rect. 5,000 18.80 31.34 1 1 3/2/5
4.1.4. Geometric Graphs
The counting task aims to predict the number of specific substructures in each graph. We collect 14 datasets (Zhao et al., 2022b). Labels include the counts of various substructures, including 3-Cycle, 4-Cycle, 5-Cycle, 6-Cycle, 7-Cycle, 8-Cycle, 4-Path, 5-Path, 6-Path, 4-Clique, Tailed-Triangle (Tailed Tri.), Chordal-Cycle, and Triangle-Rectangle.
4.2. Baselines
We compare and comprehensively investigate our proposed model with 14 representative and effective baselines in five types.
@@ 4.2.1. Node-based GNNs
  • GCN (Kipf and Welling, 2016). It utilizes graph convolutional layers to aggregate information from neighbors, serving as a fundamental GNN.
  • GIN (Xu et al., 2019). The GIN baseline employs sum pooling and utilizes Graph Isomorphism Network as the message passing layer.
  • GraphSAGE (SAGE) (Hamilton et al., 2017). It aggregates information from sampled neighbors using different functions, such as mean, LSTM, or pooling, to update node embeddings in a scalable manner.
@@ 4.2.2. Pooling-based GNNs
  • TopK (Cangea et al., 2018; Gao and Ji, 2019). It performs node drop by selecting the top-ranked nodes based on a learnable scoring function.
  • GMT (Baek et al., 2021). It condenses nodes into important ones based on grouping-matrix and create edges via attention mechanism.
  • EPool (Diehl, 2019). It progressively merges nodes by contracting high-scoring edges between adjacent nodes through edge clustering.
@@ 4.2.3. Subgraph-based GNNs
  • ECS (Yan et al., 2024). It integrates a base GNN with precomputed structural embeddings to efficiently count substructures using distance information and message passing.
  • AK+ (Zhao et al., 2022c). It extends base GNNs by replacing the local aggregation mechanism with a GNN that encodes general subgraphs instead of star-shaped patterns.
  • I2GNN (Huang et al., 2022). It enhances Subgraph MPNNs by employing multiple node identifiers, assigning distinct identifiers to the root node and its neighbors within each subgraph.
@@ 4.2.4. GL-based GNNs
  • VIBGSL (VIB) (Sun et al., 2022). It uses the Information Bottleneck principle to mask irrelevant features and learn a task-relevant graph structure for improved GNN performance.
  • HGPSL (HGP) (Zhang et al., 2019). It selects key nodes to form an induced subgraph and refines the graph structure using sparse attention-based structure learning.
  • MOSGSL (MO) (Zhou et al., 2024). It reconstructs key subgraphs using a motif-driven guidance module that captures and aligns discriminative structural patterns
@@ 4.2.5. SSL-based GNNs
  • RGC (Sun et al., 2024). It employs a diverse-curvature GCN and motif-aware contrastive learning on a stable kernel layer to model motif regularity and learn node representations.
  • MVGRL(MVG) (Hassani and Khasahmadi, 2020). It uses diffusion algorithms, such as Personalized PageRank, to generate two positive views of the graph and maximize their similarity.
  • Ours. It enhances expressivity with a $k$-path rooted subgraph approach and improves generalization through adaptive graph contrastive learning, ensuring robust performance across diverse tasks and scenarios.
4.3. Hyperparameter and Hardware Setting
For classification datasets, we set the batch size as 32 for four larger datasets (REDDIT-B, COLLAB, DD, and MolPCBA) and 128 for the other datasets. For regression datasets, we set the batch size as 256 The baselines and our model use different hyperparameters, which are fine-tuned by grid search based on validation data. For example, the hidden dimension is selected from $\{64,128,256,512\}$, the learning rate from $\{1e-2,1e-3,1e-4,1e-5\}$, and the dropout rate from $\{0,0.1,0.2,0.3,0.4,0.5\}$. Our model sets $k=1$ for graph classification, and sets $k=3$ for regression with 3-path sampling uniformly. All models are trained for a maximum of 2000 epochs, with early stopping applied if no improvement is observed on the validation set within 50 epochs.
All codes are executed on a CentOS 7 machine equipped with a 20-core Intel® Xeon® Silver 4210 CPU @ 2.20GHz, 8 NVIDIA GeForce RTX 2080 Ti GPUs, and 256GB of RAM.
@@ 4.4. Evaluation Metric
We define the effectiveness and efficiency metrics as follows.
@@ 4.4.1. Effectiveness Metric
  • Strict Accuracy (Acc). Strict accuracy is defined as $Acc=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}{\mathbb{I}(y_{i}=\hat{y}_{i})}$, where $\mathbb{I}(y_{i}=\hat{y}_{i})=1$ if only $y_{i}=\hat{y}_{i}$.
  • Micro-F1 (Mi-F1). The Micro-precision is defined as $Mi\text{-}P=\frac{\sum_{G_{i}\in\mathcal{LG}}{|y_{i}\cap\hat{y}_{i}|}}{\sum_{G_{j}\in\mathcal{LG}}{|\hat{y}_{j}|}}$ and Micro-recall is $Mi\text{-}R=\frac{\sum_{G_{i}\in\mathcal{LG}}{|y_{i}\cap\hat{y}_{i}|}}{\sum_{G_{j}\in\mathcal{LG}}{|{y}_{j}|}}$. Then, the Micro-F1 is defined as $Mi\textit{-}F1=\frac{2\times Mi\text{-}P\times Mi\text{-}R}{Mi\text{-}P+Mi\text{-}R}$.
  • Macro-F1 (Ma-F1). The Macro-precision is defined as $Ma\text{-}P=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}\frac{|y_{i}\cap\hat{y}_{i}|}{|\hat{y}_{i}|}$ and Macro-recall is defined as $Ma\text{-}R=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}\frac{|y_{i}\cap\hat{y}_{i}|}{|y_{i}|}$, Macro-F1 $Ma\text{-}F1=\frac{2\times Ma\text{-}P\times Ma\text{-}R}{Ma\text{-}P+Ma\text{-}R}$.
@@
  • Mean Absolute Error (MAE). Mean Absolute Error is defined as $MAE=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}{|y_{i}-\hat{y}_{i}|}$.
  • R2. The R2 is defined as $R2=1-\frac{\sum_{G_{i}\in\mathcal{LG}}(y_{i}-\hat{y}_{i})^{2}}{\sum_{G_{i}\in\mathcal{LG}}(y_{i}-\bar{y})^{2}}\in[0,1]$, where $\bar{y}=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}y_{i}$ is the mean of the true values. Higher $R2$ indicates better prediction performance.

### v2

@@ 4. Experiments
4.1. Evaluation Framework OpenGLT
As shown in Figure 2, we design a flexible and end-to-end evaluation framework OpenGLT that consists of three key components: the data-level, model-level, and evaluation-level components. It ensures fair and comprehensive comparisons across different methods.
  • Data-level Component. We collected 13 datasets for graph classification spanning diverse domains, including social networks, chemical graphs, and molecular graphs, as well as 13 datasets for graph regression like cycle and path counting. Additionally, we provide a unified data processing pipeline, which includes data transformation and standardized data splits tailored to various scenarios, such as imbalanced and few-shot settings. All datasets are formatted into a unified loader to facilitate fair comparisons across different methods and support future extensions.
  • Model-level Component. At the model level, we provide more than 20 GNN models to support comprehensive evaluations. These include node-based models, pooling-based models, subgraph-based models, graph-level models (GL-based), and self-supervised learning-based models (SSL-based). Additionally, we incorporate various optimization strategies tailored to GNNs, such as neighbor sampling, subgraph sampling, and graph partitioning, to enhance scalability and performance.
  • Evaluation-level Component. It focuses on ensuring fair and comprehensive comparisons of model performance. We utilize a wide range of metrics, including Macro-F1, Micro-F1, Mean Absolute Error (MAE), and $R^{2}$ scores, as well as resource-based measures such as runtime and memory consumption. To aid in the interpretation of results, we provide multiple visualization options, including tables, column charts, and line charts, enabling clear and insightful comparisons across different methods.
Our framework is designed to be user-friendly and highly extensible. Users can easily integrate new GNN models by following a standardized interface and registering them in the model library. Similarly, new datasets can be incorporated by formatting them into our unified loader, which handles preprocessing and splits automatically. The evaluation module also supports the addition of custom metrics and visualizations. To assist users, we provide comprehensive documentation with examples, ensuring that the framework can be readily extended to meet diverse research needs.
Figure 2. Evaluation framework.
@@ 4.2. Datasets
We evaluate GNNs on four domains with 26 datasets: social networks (SN), biology (BIO), chemistry (CHE), and motif counting (MC). For dataset splitting, if a standard split is publicly available and commonly used in existing works, we adopt the standard split. If there is no standard split, we use a 10-fold cross-validation approach. Detailed statistics and description are in Table 3.
@@ 4.2.1. Social Networks
  • IMDB-BINARY (I-B) (Morris et al., 2020) and IMDB-MULTI (I-M) (Morris et al., 2020). Two movie datasets from IMDB (IMDb, 2024) where each graph represents a movie, with nodes as actors and edges indicating co-appearance. The task of IMDB-B is to classify movies into two genres: Action and Romance. IMDB-M extends this to a three-class setting: Comedy, Action, and Drama.
  • REDDIT (RED) (Morris et al., 2020). We use the binary version (REDDIT-BINARY) of a dataset from Reddit (Reddit, 2024), where each graph represents a discussion thread. Nodes correspond to users, and edges indicate interactions. The task is to classify threads into two categories: Q&A-based and Discussion-based.
  • COLLAB (COL) (Morris et al., 2020). A scientific collaboration dataset where each graph represents an ego-network of a researcher, with nodes as researchers and edges indicating co-authorship. The task is to classify the networks into one of three research fields: High Energy Physics, Condensed Matter Physics, or Astrophysics.
@@ 4.2.2. Biology
  • PROTEINS (PRO) (Morris et al., 2020) and DD (Morris et al., 2020). Two binary classification datasets where each graph represents a protein. The task is to classify a protein molecule as either an enzyme or a non-enzyme. In PROTEINS, nodes represent secondary structure elements (SSEs), while in DD, nodes are amino acids; edges in both indicate spatial proximity.
  • ENZYMES (ENZ) (Morris et al., 2020). A six-class classification dataset from the BRENDA enzyme database. Each graph represents a protein, in which nodes represent SSEs with features encoding atomic properties, and edges denote spatial proximity. The task is to classify proteins into one of six enzyme commission (EC) classes.
  • MolHIV (HIV) (Hu et al., 2020a) and MolTox21 (TOX) (Hu et al., 2020a). Two molecular property prediction datasets where nodes represent atoms and edges denote chemical bonds. MolHIV is a single-label binary dataset and the task is to predict whether a molecule inhibits HIV replication. MolTox21 is a 12-label binary dataset, each label represents a specific toxicological target.
@@ 4.2.3. Chemistry
MUTAG, NCI1, MolBACE, and MolPCBA are molecular graph datasets, where nodes represent atoms and edges represent chemical bonds.
  • MUTAG (MUT) (Morris et al., 2020) and NCI1 (NCI) (Morris et al., 2020). In both datasets, node features are one-hot encodings of atomic species (e.g., C, N, O). They are binary classification tasks: MUTAG labels indicate whether a molecule is mutagenic, while NCI1 labels indicate anti-cancer activity in cell-based assays.
  • MolBACE (BAC) (Hu et al., 2020a) and MolPCBA (PCB) (Hu et al., 2020a). In both datasets, node features are derived from RDKit, including properties such as chirality and aromaticity. MolBACE is a single-label binary task for predicting whether a molecule inhibits the BACE-1 enzyme (linked to Alzheimer’s disease). MolPCBA is a 128-label binary task, each representing activity in a specific biological assay.
4.2.4. Motif-Counting Graphs
The graphs and their splits are sourced from (Chen et al., 2020) and have been widely used in follow-up works (Chen et al., 2020; Zhao et al., 2022b; Huang et al., 2022; Yan et al., 2024). Each graph is labeled with the counts of specific substructures, including Cycles (3-8 nodes), Paths (4-6 nodes), 4-Cliques, Tailed Triangles, Chordal Cycles, and Triangle-Rectangles. Graphs are generated by sampling a $d$-regular graph with $m$ nodes and randomly deleting $m$ edges, where $(m,d)$ is uniformly chosen from $\{(10,6),(15,6),(20,5),(30,5)\}$. The ground-truth counts are computed using the subgraph counting algorithm proposed in (Shervashidze et al., 2009).
Table 3. Data statistics. $\overline{\mbox{Nodes}}$ and $\overline{\mbox{Edges}}$ denote the average number of nodes and edges per graph, respectively. Split is the data partition strategy, where 10-fold denotes 10-fold cross-validation, others denote train/validation/test splits.
Type Dataset Graphs $\overline{\textbf{Nodes}}$ $\overline{\textbf{Edges}}$ Classes Split
SN IMDB-B (I-B) 1,000 19.8 96.5 2 10-fold
IMDB-M (I-M) 1,500 13.0 65.9 3 10-fold
REDDIT-B (RED) 2,000 429.6 497.8 2 10-fold
COLLAB (COL) 5,000 74.5 2457.8 3 10-fold
BIO PROTEINS (PRO) 1,113 39.1 72.8 2 10-fold
$D$ 1,178 284.3 715.7 2 10-fold
ENZYMES (ENZ) 600 32.6 62.1 6 10-fold
MolHIV (HIV) 41,127 25.5 54.9 2 8/1/1
MolTox21 (TOX) 7,831 18.9 39.2 12 8/1/1
CHE MUTAG (MUT) 188 17.9 19.8 2 10-fold
NCI1 (NCI) 4,110 29.9 32.3 2 10-fold
MolBACE (BAC) 1,513 34.1 36.9 2 8/1/1
MolPCBA (PCB) 437,929 26.0 28.1 128 8/1/1
MC {3,4,5,6,7,8}-Cycle 5,000 18.8 31.3 1 3/2/5
{4,5,6}-Path 5,000 18.8 31.3 1 3/2/5
4-Clique 5,000 18.8 31.3 1 3/2/5
Tailed Tri. 5,000 18.8 31.3 1 3/2/5
Chor. Cyc. 5,000 18.8 31.3 1 3/2/5
Tri. Rec. 5,000 18.8 31.3 1 3/2/5
4.3. Evaluated GNNs
We comprehensively evaluate 16 representative and effective GNNs across five categories as follows.
@@ 4.3.1. Node-based GNNs
  • GCN (Kipf and Welling, 2016). As a foundational GNN, GCN learns node features by aggregating information from their neighborhood, effectively performing a weighted average of neighbor representations in each layer.
  • GIN (Xu et al., 2019). GIN utilizes a Multi-Layer Perceptron (MLP) to update node representations based on a sum of its neighbors, theoretically matching the capability of the Weisfeiler-Lehman (1-WL) graph isomorphism test.
  • GraphSAGE (SAGE) (Hamilton et al., 2017). SAGE introduces a sampling-based neighborhood aggregation strategy. For each node, it uniformly samples a fixed number of neighbors to generate embeddings, making the learning process scalable and inductive.
  • PNA (Corso et al., 2020). PNA combines multiple aggregators (e.g., mean, max, sum) with degree-based scalers to capture the distribution of neighbor features, allowing the model to distinguish diverse neighbor patterns better than single-aggregator methods.
@@ 4.3.2. Pooling-based GNNs
  • TopK (Cangea et al., 2018; Gao and Ji, 2019). TopK uses a trainable projection vector to score nodes, with the top-k nodes passing GNN backbones selected to form the graph for the subsequent layer.
  • GMT (Baek et al., 2021). GMT introduces a hierarchical pooling method where nodes are adaptively grouped into clusters using a transformer-based attention mechanism, allowing the model to learn soft cluster assignments and generate coarse-grained graph representations at multiple scales.
  • EdgePool (EPool) (Diehl, 2019). EPool learns a hierarchical representation by iteratively contracting edges. It scores edges based on their importance for connecting nodes and merges the corresponding node pairs, effectively simplifying the graph topology layer by layer.
@@ 4.3.3. Subgraph-based GNNs
  • ECS (Yan et al., 2024). ECS boosts the GNN backones by injecting pre-calculated structural information. It computes embeddings that encode subgraph-centric distances, allowing the model to count substructures with provable guarantees.
  • GNNAK+ (AK+) (Zhao et al., 2022a). AK+ shifts the computational unit from single-node neighborhoods to k-hop subgraphs, thereby enabling the GNN backbones to capture more complex, high-order structural patterns.
  • I2GNN (I2) (Huang et al., 2022). I2GNN breaks structural symmetries by assigning unique identifiers to the central root node and its neighbors within each sampled subgraph, enabling finer-grained representation learning.
@@ 4.3.4. GL-based GNNs
  • VIBGSL (VIB) (Sun et al., 2022). VIB uses the Information Bottleneck principle to mask irrelevant features and learn a task-relevant graph structure for better modeling the key patterns.
  • HGP-SL (HGP) (Zhang et al., 2019). HGP selects key nodes to form an induced subgraph and refines the graph structure using sparse attention-based structure learning.
  • MOSGSL (MO) (Zhou et al., 2024). MO reconstructs key subgraphs using a motif-driven guidance module that dynamically captures and aligns discriminative structural patterns.
@@ 4.3.5. SSL-based GNNs
  • RGC (Sun et al., 2024). RGC employs a diverse-curvature GCN and motif-aware contrastive learning on a stable kernel layer to robustly model motif regularity and learn node representations.
  • MVGRL (MVG) (Hassani and Khasahmadi, 2020). MVG learns graph representations by contrasting node embeddings from the original and diffusion-based views of the graph.
  • GCA (Zhu and others, 2021). GCA learns graph representations by contrasting node embeddings across adaptively augmented views that preserve important structural and semantic information.
@@ 4.4. Evaluation Metric
We evaluate the performance of GNNs using effectiveness and efficiency metrics as follows.
@@ 4.4.1. Effectiveness Metric
  • Strict Accuracy (Acc). Strict accuracy is used to measure the proportion of exact matches between predicted and true labels. Strict accuracy is defined as $Acc=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}{\mathbb{I}(y_{i}=\hat{y}_{i})}$, where $\mathbb{I}(y_{i}=\hat{y}_{i})=1$ if only $y_{i}=\hat{y}_{i}$.
  • Micro-F1 (Mi-F1). Micro-F1 is a performance metric that considers the overall precision and recall across all instances in the dataset. The Micro-precision is defined as $Mi\text{-}P=\frac{\sum_{G_{i}\in\mathcal{LG}}{|y_{i}\cap\hat{y}_{i}|}}{\sum_{G_{j}\in\mathcal{LG}}{|\hat{y}_{j}|}}$ and Micro-recall is $Mi\text{-}R=\frac{\sum_{G_{i}\in\mathcal{LG}}{|y_{i}\cap\hat{y}_{i}|}}{\sum_{G_{j}\in\mathcal{LG}}{|{y}_{j}|}}$. Then, the Micro-F1 is defined as $Mi\textit{-}F1=\frac{2\times Mi\text{-}P\times Mi\text{-}R}{Mi\text{-}P+Mi\text{-}R}$.
  • Macro-F1 (Ma-F1). Macro-F1 evaluates the average performance of precision and recall across all instances, treating each equally regardless of size. The Macro-precision is defined as $Ma\text{-}P=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}\frac{|y_{i}\cap\hat{y}_{i}|}{|\hat{y}_{i}|}$ and Macro-recall is defined as $Ma\text{-}R=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}\frac{|y_{i}\cap\hat{y}_{i}|}{|y_{i}|}$, so the Macro-F1 is defined as $Ma\text{-}F1=\frac{2\times Ma\text{-}P\times Ma\text{-}R}{Ma\text{-}P+Ma\text{-}R}$.
@@
  • Mean Absolute Error (MAE). Mean Absolute Error measures the average magnitude of errors between predicted and true values. MAE is defined as $MAE=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}{|y_{i}-\hat{y}_{i}|}$. Lower $MAE$ indicates better performance.
  • R2. R2 evaluates the proportion of variance in the true values that is captured by the predicted values. The R2 is defined as $R2=1-\frac{\sum_{G_{i}\in\mathcal{LG}}(y_{i}-\hat{y}_{i})^{2}}{\sum_{G_{i}\in\mathcal{LG}}(y_{i}-\bar{y})^{2}}\in[0,1]$, where $\bar{y}=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}y_{i}$ is the mean of the true values. Higher $R2$ indicates better performance.